### v1

# Protect Measurement-Induced Phase Transition from Noise

Abstract
Measurement-induced phase transition (MIPT) is a novel non-equilibrium phase transition characterized by entanglement entropy. The scrambling dynamics induced by random unitary gates can protect information from low-rate measurements. However, common decoherence noises, such as dephasing, are detrimental to the volume law phase, posing a significant challenge for observing MIPT in current noisy intermediate-scale quantum devices. Here, we demonstrate that incorporating quantum-enhanced operations can effectively protect MIPT from environmental noise. The conditional entanglement entropy is associated with a statistical mechanics model wherein noise and quantum-enhanced operations act as two competing external random fields. Then we show that an average apparatus-environment exchange symmetry ensures the conditional entanglement entropy is a valid probe of entanglement. Furthermore, we provide numerical evidence on a (2+1)-d quantum circuit under dephasing noise, demonstrating that MIPT can indeed be observed with the aid of quantum-enhanced operations. This result not only serves as a concrete example of the power of quantum enhancement in combating noise but also holds experimental relevance, as the protocol is straightforward to implement in practice.
Noise presents one of the most significant threats to reliable quantum computation and the long-term storage of quantum information. To address this challenge, quantum error correction (QEC) techniques are employed [1, 2, 3]. The core idea of QEC is to encode quantum information in a noise-resilient manner, allowing errors induced by noise to be effectively detected and corrected. The encoding schemes and correction procedures in QEC are typically highly structured [4, 5].
In recent years, measurement-induced phase transitions (MIPT) have attracted significant attention [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. MIPTs can be understood from the perspective of QEC, where information is encoded in an unstructured way through scrambling by random unitary gates, with measurements identified as sources of noise [30]. When the measurement rate is low, corresponding to a low error rate, the entanglement entropy of the state follows a volume law, indicating that the information remains protected. Although this transition has been demonstrated in several experiments [31, 32, 33, 34], a major obstacle to observing this transition is its instability against various decoherence noises. It has been shown that the state obeys an area law even with an infinitesimal rate of dephasing noise or resetting noise in the circuit’s bulk [35, 36, 37, 38]. This can be interpreted as the scrambling dynamics being insufficient to protect quantum information against these more common practical noises. Therefore, it is both experimentally relevant and theoretically interesting to find a way to protect MIPT from decoherence noise.
Meanwhile, quantum-enhanced (QE) operation has emerged as a potentially more powerful and flexible method for extracting information about a quantum state compared to traditional projective measurements, which only allow access to classical information [39, 40, 41, 42]. The fundamental idea of quantum enhancement is to use a quantum sensor, rather than a classical sensor, to detect the system. It has been demonstrated that an exponential speedup can be achieved in certain tasks by coherently manipulating these quantum probes [39]. However, it remains an open question whether these QE operations and algorithms can maintain their advantage in the presence of noise.
In this Letter, we demonstrate that the original MIPT can be protected against decoherence caused by environmental noise through QE operations. From the perspective of QEC, these QE operations act as an unstructured protocol to protect quantum information by encoding it within both the system qubits and the ancilla qubits.
We first provide analytical analysis, mapping the random quantum circuit to a statistical-mechanics model. We find that noise and QE operation exactly map to two symmetry-breaking fields in different directions. Combining with unitary gate and measurement, we obtain a random field model with ferromagnetic coupling that supports a ferromagnetic-paramagnetic phase transition with both increasing temperature and increasing random field strength. Next, we discover that an average apparatus-environment exchange (aAEE) symmetry ensures the conditional entanglement entropy (CEE) to be a valid probe of entanglement [43, 44]. By imposing the aAEE symmetry, we conduct numerical studies for (2+1)-d circuits and explicitly demonstrate the MIPT in the presence of various decoherence noises. Our approach is inspired by Ref. [43], where a prototype setting was introduced. The structure we introduced here incorporates measurement and is much easier to implement in practice, as will be further elaborated later.
Figure 1: Circuit structure and operations. (a) The quantum circuit model we consider in this work. Orange, green, red, and blue rectangles represent random unitary gates, measurements, noise, and QE operations, respectively. Noise and QE operation are further detailed as interactions with environment and ancilla qubits. (b) Dephasing noise and the corresponding symmetric QE operation. $U_{SE}=U_{SA}=\text{CNOT}$ and $\rho_{E}=\rho_{A}=\left|0\right\rangle\left\langle 0\right|$. (c) Depolarizing noise and the corresponding symmetric QE operation. Notice that it requires two environment qubits to represent the noise. $U_{SE}=U_{SA}=\text{SWAP}$. $\rho_{E}$ and $\rho_{A}$ are in two-qubit Bell state.
Circuit model. We consider a quantum circuit where four types of operations are allowed: random unitary gate, measurement, noise, and QE operation. As shown in Fig. 1(a), the unitary gates are applied in a brick wall pattern. Between every two unitary layers, each qubit has a probability $p$ of being measured projectively. Meanwhile, noise or QE operations occur with probabilities $q_{n}$ and $q_{e}$, respectively. Although noises are usually represented as quantum channels, we can adopt the Stinespring representation of the channel to represent it by unitary operations. Specifically, a quantum channel acting on the system can be written as $\mathcal{N}(\rho)=\text{Tr}_{E}(U_{SE}(\rho\otimes\rho_{E})U_{SE}^{\dagger})$ where $U_{SE}$ represents the interaction between the system and environment, and $\rho_{E}$ is the initial state of the environment [45]. In the left column of Fig. 1(b) and 1(c), we show two examples of representing dephasing and depolarizing noise from this perspective. Similarly, QE operations can be considered as applying a unitary $U_{SA}$ to entangle the system with the ancilla qubits, whose initial states are $\rho_{A}$. We provide representative examples in the right column of Fig. 1(b) and 1(c). It’s worth noticing that the number of environment and ancilla qubits are not limited. Despite the similarity in the diagrams, it is crucial to emphasize that the environment qubits are always discarded at the end, signifying that the information is inevitably lost by tracing out these qubits. In contrast, the ancilla qubits remain under our control, allowing us to have full access to and manipulation of them.
The input global state can be written as $\left|\Psi_{0}^{ESA}\right\rangle=\left|\psi_{E}\right\rangle\otimes\left|\psi_{S}\right\rangle\otimes\left|\psi_{A}\right\rangle$ and a certain realization can be regarded as acting unitary gates and projection operators on it. At the end of every realization ${m}$, the global state can be represented as an unnormalized pure state $\left|\Psi_{m}^{ESA}\right\rangle$ and the physically relevant state is $\rho_{m}^{SA}=\text{Tr}_{E}\left(\left|\Psi_{m}^{ESA}\right\rangle\left\langle\Psi_{m}^{ESA}\right|\right)$. For convenience, we drop the superscript from now on. In this work, the realization $m$ includes both the circuit structure resulting from the random gates and operation locations, as well as the trajectory labeling by different measurement outcomes.
Figure 2: Statistical-mechanics model. (a) $\overline{S(M|A)}$ is equal to the free energy difference of the same random field model under different boundary conditions. Red and blue color represents $\mathbb{I}$ and $\mathbb{C}$, respectively. The circles within the bulk indicate that the boundary conditions on environment and ancilla qubits are equivalent to applying extrinsic fields in different directions on the spins. Their locations are random due to the stochastic application of noise and QE operations. (b), (c) Detailed schematics illustrating the bond weight in the statistical-mechanics model, showing a qubit subjected to resetting noise and QE operations.
Analytical analysis. Instead of considering subsystem entanglement entropy, we aim to calculate the CEE $\overline{S(M|A)}\equiv\overline{S(M,A)}-\overline{S(A)}$. The overline here represents averaging all circuits and measurement outcomes. Mathematically, this can be expressed as,
$\overline{S(M|A)}$ $=$ $\lim_{n\to 1}(\overline{S^{(n)}(M,A)}-\overline{S^{(n)}(A)})$ (1)
$=$ $\lim_{n\to 1}\lim_{k\to 0}\frac{1}{(1-n)k}\text{log}\left(\frac{\mathcal{Z}^{(n,k)}_{MA}}{\mathcal{Z}^{(n,k)}_{A}}\right)$
$=$ $\lim_{n\to 1}\lim_{k\to 0}\frac{1}{(n-1)k}\left(\mathcal{F}^{(n,k)}_{MA}-\mathcal{F}^{(n,k)}_{A}\right),$
where
$\mathcal{Z}_{MA}^{(n,k)}$ $=$ $\sum_{m}\text{Tr}\left(\mathbb{C}^{M}\Lambda_{A}^{(n,k)}\right),\mathcal{Z}_{A}^{(n,k)}=\sum_{m}\text{Tr}\left(\Lambda_{A}^{(n,k)}\right),$
$\Lambda_{A}^{(n,k)}$ $=$ $\text{Tr}_{A\cup E}\left[\left(\left|\Psi_{m}\right\rangle\left\langle\Psi_{m}\right|\right)^{\otimes Q}\mathbb{C}^{A}\right].$ (2)
$\mathcal{F}^{n,k}=-\text{log}\mathcal{Z}^{(n,k)}$ is identified as the free energy [46]. The underlying statistical-mechanics model for $\mathcal{Z}^{(n,k)}$ is composed of spins that take values in the permutation group $\mathbb{S}(Q)$ with $Q=nk+1$. We designate the identity operation in $\mathbb{S}(Q)$ as $\mathbb{I}_{Q}$ and a particular permutation operator as $\mathbb{C}=\mathbb{S}^{\otimes k}\otimes\mathbb{I}$. $\mathbb{S}$ is the cyclic permutation operation in permutation group $\mathbb{S}(n)$. Thus, $\overline{S(M|A)}$ is related to the free energy difference under different boundary conditions, as shown in Fig. 2(a). In both $\mathcal{Z}_{MA}$ and $\mathcal{Z}_{A}$, the environment qubits are always subject to the boundary condition $\mathbb{I}$ while the ancilla qubits are subject to $\mathbb{C}$. It has been shown that the entanglement phase transition can be seen as the spontaneous breaking of the $(\mathbb{S}_{Q}\times\mathbb{S}_{Q})\rtimes\mathbb{Z}_{2}$ symmetry [47, 48, 49]. However, boundary condition imposing on environment qubits can be effectively understood as an explicit symmetry-breaking field in the bulk that align the spins to direction $\mathbb{I}$, which is thus detrimental to MIPT [35, 36, 37]. Through QE operation and considering the conditional entropy, we effectively introduced another symmetry-breaking field in direction $\mathbb{C}$ to counteract with the effect of noises. Taking the example of the resetting noise and its associated symmetric QE operation $U_{SE}=U_{SA}=\text{SWAP},\rho_{E}=\rho_{A}=\left|0\right\rangle\left\langle 0\right|$, the bond weight connecting two spins $\sigma$ and $\tau$ can be explicitly formulated as [46],
$W^{\text{reset}}(\sigma,\tau)$ $=$ $(1-p)(1-q)\left\langle\sigma|\tau\right\rangle$ (3)
$+(1-p)\frac{q}{2}\left(\left\langle\sigma|\mathbb{C}\right\rangle+\left\langle\sigma|\mathbb{I}\right\rangle\right)+pd.$
We choose $q_{n}=q_{e}=q/2$ which ensures that the total symmetry-breaking field is zero. The model has a $(\mathcal{C}_{\mathbb{S}_{Q}}(\mathbb{C})\rtimes\mathbb{Z}_{2})\rtimes\mathbb{Z}_{2}$ symmetry where $\mathcal{C}_{\mathbb{S_{Q}}}(\mathbb{C})$ is the centralizer of $\mathbb{C}$ [46]. Note that $\left\langle\sigma|\tau\right\rangle=d^{|\sigma\tau^{-1}|}$ where $d$ is the local Hilbert space dimension and $|\sigma\tau^{-1}|$ denotes the number of cycles in permutation $\sigma\tau^{-1}$. In the $d\to\infty$ limit, $\left\langle\sigma|\tau\right\rangle$ becomes $d^{Q}\delta_{\sigma,\tau}$. This corresponds to a random field $Q!$-state Potts model, where the first term can be interpreted as a ferromagnetic coupling, the second term as two competing random fields, and the final term as the effect of temperature. The strength of the random fields is controlled by $q$ while temperature is controlled by $p$. Given prior work demonstrating phase transitions in random field Ising models and Potts models [50, 51, 52, 53, 54, 55, 56, 57, 58, 59], we anticipate that a phase transition would occur at finite $p$ and $q$ in quantum circuit which is directly characterized by $\overline{S(M|A)}$. For depolarizing and dephasing noises, the corresponding statistical-mechanics models are analogous as provided in Supplemental Material [46].
Figure 3: Relationship between symmetries. (a) AEE symmetry requires that every realization remains invariant under the SWAP operation on environment and ancilla qubits, which consequently leads to IE symmetry. (b) The aAEE symmetry is the averaged version of AEE symmetry. The dashed lines indicate that unentangled qubits can be added to balance the number of qubits before applying the SWAP operation. The aAEE symmetry would then result in an aIE symmetry, which is the averaged version of IE symmetry.
Average symmetry. Since $\overline{S(M|A)}$, as opposed to $\overline{S(M)}$ in traditional setting, undergoes a phase transition, it is pertinent to determine under what conditions $\overline{S(M|A)}$ serves as a valid probe of entanglement or quantum information [60]. It has been proven that if the system has information exchange (IE) symmetry in every realization, such that $S(M|A)=S(M|E)$, then $S(M|A)$ ensures entanglement between $M$ and $M^{c}\cup A$ [44]. Here, $M^{c}$ denotes the complement of the subsystem $M$. Volume law scaling of $S(M|A)$ would then yield an extensive number of distillable Bell pairs between $M$ and $M^{c}\cup A$. The IE symmetry can be guaranteed if an apparatus-environment exchange (AEE) symmetry is present, which requires that each circuit realization be symmetric under the exchange of ancilla and environment qubits. However, this symmetry is explicitly violated by any realization in our circuit since noise and QE operation are independently chosen and applied at each location. Additionally, the number of ancilla and environment qubits may differ within a specific circuit realization, making the “exchange” operation more subtle to define. To solve this issue, we propose that an aAEE symmetry can still lead to an average IE (aIE) symmetry, thereby making the averaged CEE a valid probe of entanglement. The relationship between these symmetries are summarized in Fig. 3. We first define a generalized SWAP operation $\text{SWAP}^{\prime}$ on $\left|\Psi_{m}\right\rangle$ as initially adding unentangled qubits to either $E$ or $A$ to equalize the number of qubits. Subsequently, a conventional SWAP gate is applied to $E$ and $A$, followed by the removal of the unentangled qubits. $\text{SWAP}^{\prime}\left|\Psi_{m}\right\rangle$ would not necessarily equal to $\left|\Psi_{m}\right\rangle$, but may transform it into another realization $\left|\Psi_{m^{\prime}}\right\rangle$. The aAEE symmetry requires that the probabilities of these two realizations occurring are identical. Concretely, one can represent the circuit’s outcome as $\rho=\sum_{m}\left|\Psi_{m}\right\rangle\left\langle\Psi_{m}\right|$, and the aAEE symmetry necessitates that
$\text{SWAP}^{\prime}\rho\text{SWAP}^{\prime\dagger}=\rho.$ (4)
This is reminiscent of the recently proposed average (weak) symmetry in the study of mixed state order [61, 62, 63, 64, 65, 66]. A subtlety in defining $\rho$ in this manner is that the number of qubits in $\left|\Psi_{m}\right\rangle$ may vary, but this can be simply addressed by adding unentangled qubits to states with fewer qubits. With aAEE symmetry and combining with Eq. (2), one would have
$\sum_{m}\Lambda_{A}^{Q}$ $=$ $\text{Tr}_{A\cup E}(\rho^{\otimes Q}\mathbb{C}^{A})=\text{Tr}_{A\cup E}\left(\text{SWAP}^{\prime}\rho^{\otimes Q}\text{SWAP}^{\prime\dagger}\right.$ (5)
$\left.\mathbb{C}^{A}\right)=\text{Tr}_{A\cup E}(\rho^{\otimes Q}\mathbb{C}^{E})\equiv\sum_{m}\Lambda^{Q}_{E}.$
This result implies that $\overline{S(M|A)}=\overline{S(M|E)}$, which we call the aIE symmetry. It is then straightforward to show that aIE symmetry results in,
$\overline{S(M|A)}\geq 0,\quad\overline{S(M|A)}=\overline{S(M^{c}|A)},$
$\overline{S(M|A)}=-\overline{S(M|A\cup M^{c})},$ (6)
which also ensure that $\overline{S(M|A)}$ represents the entanglement between $M$ and $M^{c}\cup A$.
Translating the aAEE symmetry into the concrete circuit model reveals that it requires $\rho_{E}=\rho_{A}$, $U_{SE}=U_{SA}$ and $q_{n}=q_{e}=q/2$. Thus, QE operation should be symmetric with the corresponding noise, and the operation should be applied at the same rate as the noise. It is noteworthy that aAEE symmetry condition $q_{n}=q_{e}$ aligns with the condition of zero net field in the statistical-mechanics model. Since every noise can be represented in the Stinespring form [45], QE operations can protect MIPT from various types of noise.
Figure 4: Numerical results for dephasing noise. (a) We partition the system into four neighboring regions with equal size as a, b, c, d. The conditional tripartite mutual information is then calculated as $I_{3}=S^{\prime}(\text{a})+S^{\prime}(\text{b})+S^{\prime}(\text{c})-S^{\prime}(\text{ab})-S^{\prime}(\text{ac})-S^{\prime}(\text{bc})+S^{\prime}(\text{abc})$ where $S^{\prime}(x)\equiv\overline{S(x|A)}$. (b) $I_{3}$ vs $(p,q)$ with $L=20$. The red line $q=0.1$ is further examined in (c), (d), which demonstrate MIPT in the presence of noise and QE operation. Each data point is averaged over $5\times 10^{4}$ realizations.
Numerical results. We now turn to numerical calculations to explicitly demonstrate MIPT in the presence of noise and QE operations. Considering the absence of phase transitions in the 2D random field Ising model due to Imry-Ma argument [58], we directly simulate the (2+1)-d quantum circuit with $L\times L$ qubits on a square lattice, corresponding to a 3D statistical model. To ensure efficient simulation, we choose the random unitary gates to be Clifford gates, allowing the circuit to be simulated using the stabilizer formalism [67, 68, 69]. A single time step consists of four random unitary layers [46]. Between every two layers, measurement in $z$-basis, noise, and QE operations occur at each site with probabilities $p$, $q/2$ and $q/2$, respectively. In the absence of any noise, MIPT occurs at around $p\sim 0.3$. We evolve the circuit for depth $T=10L$ to achieve convergence in all simulations and employ the technique proposed in Ref. [43] to perform efficient simulations without storing the ancilla qubits explicitly. In particular, we consider dephasing noise here, and results for other types of noise and additional numerical results can be found in Supplemental Material [46].
To determine the critical point, we consider the conditional tripartite mutual information $I_{3}$. We take periodic boundary condition in both directions and partition the system into four segments with equal size, as depicted in Fig. 4(a). $I_{3}$ is expected to scale with $L$ in the volume law phase, while approach 00 in the area law phase [70, 71]. The complete phase diagram can thus be constructed by identifying the regions where $I_{3}$ is zero or not, as shown in Fig. 4(b), where a volume law phase is clearly observed for non-zero values of $q$ and $p$. We further focus on the line $q=0.1$ and collapse the data according to the scaling form $I_{3}=f((p-p_{c})L^{1/\nu})$. The results are shown in Fig. 4(c) and 4(d). The critical point is $p_{c}=0.214(2)$ with the critical exponent $\nu=0.9(1)$, which is close to the result for noiseless case in [71]. Details for data collapse are provided in [46].
Discussions. Similar to QEC, where more physical qubits are required to encode a smaller number of logical qubits, our scheme necessitates the preparation of additional ancilla qubits as the circuit depth increases. Specifically, $O(qNT)$ ancilla qubits are needed, where $N$ is the number of system qubits, $p$ is the noise rate, and $T$ is the circuit depth. Assuming $q_{e}\sim 1\%,L=20$ and $T=10L$, around $800$ ancilla qubits are needed to observe the phase transition. It is important to highlight that each ancilla qubit is utilized only once and can be isolated after use, thereby mitigating the risk of noise contamination.
An immediate application of our scheme requires prior knowledge of the probability that a specific type of decoherence noise will occur in the experimental setting. Our scheme can be directly adapted to determine the noise rate as follows: apply a finite amount of measurement such that the state enters the area-law phase, then sweep through different QE operation rates $q_{e}$. In a finite-size system, $\overline{S(M|A)}$ will be greater than zero if $q_{e}>q_{n}$ and less than zero otherwise.
Besides noise, post-selection is another significant challenge preventing MIPT from being easily observed. Here the post-selection problem persists when the measurement rate is non-zero because determining the CEE of a single trajectory requires post-selecting on the measurement outcomes, a difficulty rooted in Born’s rule. Recent work has utilized various post-selection-free methods to probe the phase transition, such as using cross-entropy benchmarking [72] or providing bounds on the entanglement entropy [73, 74]. We believe these methods can also be extended to the conditional entropy in this work, and we leave this exploration for future work.
Finally, we highlight the differences between our circuit model and that in Ref. [43, 44]. Apart from incorporating measurement in our circuit, the most notable difference is that noise and QE operations act randomly and independently, not necessarily at the same location. The only requirement is that their rates are equal, which serves as a more realistic and feasible setting since we usually cannot accurately predict where and when noise will occur. Additionally, previous work has demonstrated a noise-induced phase transition where the noise rate is inversely proportional to the system size [75]. In our case, the noise rate is constant, which again represents a more realistic scenario.
Acknowledgements.
Acknowledgment. We thank Shane P. Kelly for valuable discussions. This work is supported by the National Key Research Program of China under Grant No. 2019YFA0308404, the Natural Science Foundation of China through Grants No. 12350404 and No. 12174066, the Innovation Program for Quantum Science and Technology through Grant No. 2021ZD0302600, the Science and Technology Commission of Shanghai Municipality under Grants No. 23JC1400600 and No. 2019SHZDZX01.

### v2

@@
 (2), one would have
-$\sum_{m}\Lambda_{A}^{Q}$ $=$ $\text{Tr}_{A\cup E}(\rho^{\otimes Q}\mathbb{C}^{A})=\text{Tr}_{A\cup E}\left(\text{SWAP}^{\prime}\rho^{\otimes Q}\text{SWAP}^{\prime\dagger}\right.$ (5)
-$\left.\mathbb{C}^{A}\right)=\text{Tr}_{A\cup E}(\rho^{\otimes Q}\mathbb{C}^{E})\equiv\sum_{m}\Lambda^{Q}_{E}.$
+$\sum_{m}\Lambda_{A}^{Q}$ $=\text{Tr}_{A\cup E}(\rho^{\otimes Q}\mathbb{C}^{A})=\text{Tr}_{A\cup E}\left(\text{SWAP}^{\prime}\rho^{\otimes Q}\right.$ (5)
+$\left.\text{SWAP}^{\prime\dagger}\mathbb{C}^{A}\right)=\text{Tr}_{A\cup E}(\rho^{\otimes Q}\mathbb{C}^{E})\equiv\sum_{m}\Lambda^{Q}_{E}.$
 This result implies that $\overline{S(M|A)}=\overline{S(M|E)}$, which we call the aIE symmetry. It is then straightforward to show that aIE symmetry results in,
@@
 Assuming $q_{e}\sim 1\%,L=20$ and $T=10L$, around $800$ ancilla qubits are needed to observe the phase transition.
-It is important to highlight that each ancilla qubit is utilized only once and can be isolated after use, thereby mitigating the risk of noise contamination.
+It is imperative to underscore that we ignore potential noise in the ancilla qubits.
+This is justified, as each ancilla qubit is utilized only once and can be effectively isolated after use, thereby mitigating the risk of noise propagation and contamination.
 An immediate application of our scheme requires prior knowledge of the probability that a specific type of decoherence noise will occur in the experimental setting.
 Our scheme can be directly adapted to determine the noise rate as follows: apply a finite amount of measurement such that the state enters the area-law phase, then sweep through different QE operation rates $q_{e}$.
-In a finite-size system, $\overline{S(M|A)}$ will be greater than zero if $q_{e}>q_{n}$ and less than zero otherwise.
+In the thermodynamic limit, curves corresponding to different $p$ would intersect at the point $q_{n}=q_{e}$.
+A more detailed explanation and finite size numerical results are provided in [46].
 Besides noise, post-selection is another significant challenge preventing MIPT from being easily observed.

### v3

@@
 Abstract
-Measurement-induced phase transition (MIPT) is a novel non-equilibrium phase transition characterized by entanglement entropy.
-The scrambling dynamics induced by random unitary gates can protect information from low-rate measurements.
-However, common decoherence noises, such as dephasing, are detrimental to the volume law phase, posing a significant challenge for observing MIPT in current noisy intermediate-scale quantum devices.
-Here, we demonstrate that incorporating quantum-enhanced operations can effectively protect MIPT from environmental noise.
-The conditional entanglement entropy is associated with a statistical mechanics model wherein noise and quantum-enhanced operations act as two competing external random fields.
-Then we show that an average apparatus-environment exchange symmetry ensures the conditional entanglement entropy is a valid probe of entanglement.
-Furthermore, we provide numerical evidence on a (2+1)-d quantum circuit under dephasing noise, demonstrating that MIPT can indeed be observed with the aid of quantum-enhanced operations.
-This result not only serves as a concrete example of the power of quantum enhancement in combating noise but also holds experimental relevance, as the protocol is straightforward to implement in practice.
+Scrambling dynamics induced by random unitary gates can protect information from low-rate measurements, which underpins the phenomenon known as the measurement-induced phase transition (MIPT).
+However, typical decoherence noises disrupts the volume law phase, complicating the observation of MIPT on noisy intermediate-scale quantum devices.
+Here, we demonstrate that incorporating quantum-enhanced operations can effectively protect MIPT from environmental noise, thereby enabling its detection in experiment.
+The transition is characterized by the conditional entanglement entropy (CEE), which is associated with a statistical mechanics model wherein noise and quantum-enhanced operations act as competing external random fields.
+When the net external field is zero, a ferromagnetic-paramagnetic phase transition is expected, resulting in the MIPT.
+This zero-field condition also ensures an average apparatus-environment symmetry, making CEE a valid probe of entanglement and establishing the transition as a genuine entanglement phase transition.
+Additionally, we provide numerical results demonstrate the MIPT in a (2+1)-dimensional quantum circuit under dephasing noise.
+We also propose a method to estimate the noise rate, enabling the zero-field condition to be achieved experimentally and ensuring the feasibility of our protocol.
+Our result serves as a concrete example of the power of quantum enhancement in combating noise.
 Noise presents one of the most significant threats to reliable quantum computation and the long-term storage of quantum information.
@@
 From the perspective of QEC, these QE operations act as an unstructured protocol to protect quantum information by encoding it within both the system qubits and the ancilla qubits.
-We first provide analytical analysis, mapping the random quantum circuit to a statistical-mechanics model.
-We find that noise and QE operation exactly map to two symmetry-breaking fields in different directions.
-Combining with unitary gate and measurement, we obtain a random field model with ferromagnetic coupling that supports a ferromagnetic-paramagnetic phase transition with both increasing temperature and increasing random field strength.
-Next, we discover that an average apparatus-environment exchange (aAEE) symmetry ensures the conditional entanglement entropy (CEE) to be a valid probe of entanglement [43, 44].
+We first provide analytical analysis, mapping the conditional entanglement entropy (CEE) to the free energy of a statistical-mechanics model, where noise and QE operation exactly correspond to two symmetry-breaking fields in different directions.
+Combining with unitary gate and measurement, we obtain a random field model with ferromagnetic coupling.
+This model exhibits a ferromagnetic-paramagnetic phase transition driven by both increasing temperature and increasing random field strength, under the condition of zero net external field.
+This phase transition thus leads to the transition in the scaling of CEE.
+Next, we discover that the zero-field condition is equivalent to an average apparatus-environment exchange (aAEE) symmetry, which ensures the CEE to be a valid probe of entanglement [43, 44].
+Consequently, the transition observed in CEE directly corresponds to an entanglement phase transition.
 By imposing the aAEE symmetry, we conduct numerical studies for (2+1)-d circuits and explicitly demonstrate the MIPT in the presence of various decoherence noises.
@@
 Instead of considering subsystem entanglement entropy, we aim to calculate the CEE $\overline{S(M|A)}\equiv\overline{S(M,A)}-\overline{S(A)}$.
-The overline here represents averaging all circuits and measurement outcomes.
+$M$ denotes an arbitrary subsystem and the overline here represents averaging all circuits and measurement outcomes.
 Mathematically, this can be expressed as,
@@
 $\mathcal{F}^{n,k}=-\text{log}\mathcal{Z}^{(n,k)}$ is identified as the free energy [46].
-The underlying statistical-mechanics model for $\mathcal{Z}^{(n,k)}$ is composed of spins that take values in the permutation group $\mathbb{S}(Q)$ with $Q=nk+1$.
-We designate the identity operation in $\mathbb{S}(Q)$ as $\mathbb{I}_{Q}$ and a particular permutation operator as $\mathbb{C}=\mathbb{S}^{\otimes k}\otimes\mathbb{I}$.
-$\mathbb{S}$ is the cyclic permutation operation in permutation group $\mathbb{S}(n)$.
-Thus, $\overline{S(M|A)}$ is related to the free energy difference under different boundary conditions, as shown in Fig.
+The underlying statistical-mechanics model for $\mathcal{Z}^{(n,k)}$ consists of spins that take values in the permutation group $\mathbb{S}(Q)$ with $Q=nk+1$, where $\mathbb{I}$ and $\mathbb{C}$ correspond to two particular group element in $\mathbb{S}(Q)$ [46].
+A critical observation is that, in both $\mathcal{Z}_{MA}$ and $\mathcal{Z}_{A}$, the environment qubits are subject to the same boundary condition $\mathbb{I}$, while the ancilla qubits are subject to $\mathbb{C}$.
+Consequently, we can interpret them as competing external fields applied at random positions within the bulk, with $\Lambda_{A}^{(n,k)}$ representing the bulk partition function.
+In other words, through QE operation and considering the conditional entropy, a symmetry-breaking field along the $\mathbb{C}$ direction is effectively introduced to counteract the influence of noise.
+Moreover, the unitary gates induce a ferromagnetic coupling, while the measurement rate acting as the temperature [46].
+Thus, $\overline{S(M|A)}$ is directly related to the free energy difference of a random field model with ferromagnetic coupling, subject to different boundary conditions, as shown in Fig.
 2(a).
-In both $\mathcal{Z}_{MA}$ and $\mathcal{Z}_{A}$, the environment qubits are always subject to the boundary condition $\mathbb{I}$ while the ancilla qubits are subject to $\mathbb{C}$.
-It has been shown that the entanglement phase transition can be seen as the spontaneous breaking of the $(\mathbb{S}_{Q}\times\mathbb{S}_{Q})\rtimes\mathbb{Z}_{2}$ symmetry [47, 48, 49].
-However, boundary condition imposing on environment qubits can be effectively understood as an explicit symmetry-breaking field in the bulk that align the spins to direction $\mathbb{I}$, which is thus detrimental to MIPT [35, 36, 37].
-Through QE operation and considering the conditional entropy, we effectively introduced another symmetry-breaking field in direction $\mathbb{C}$ to counteract with the effect of noises.
-Taking the example of the resetting noise and its associated symmetric QE operation $U_{SE}=U_{SA}=\text{SWAP},\rho_{E}=\rho_{A}=\left|0\right\rangle\left\langle 0\right|$, the bond weight connecting two spins $\sigma$ and $\tau$ can be explicitly formulated as [46],
-$W^{\text{reset}}(\sigma,\tau)$ $=$ $(1-p)(1-q)\left\langle\sigma|\tau\right\rangle$ (3)
-$+(1-p)\frac{q}{2}\left(\left\langle\sigma|\mathbb{C}\right\rangle+\left\langle\sigma|\mathbb{I}\right\rangle\right)+pd.$
-We choose $q_{n}=q_{e}=q/2$ which ensures that the total symmetry-breaking field is zero.
-The model has a $(\mathcal{C}_{\mathbb{S}_{Q}}(\mathbb{C})\rtimes\mathbb{Z}_{2})\rtimes\mathbb{Z}_{2}$ symmetry where $\mathcal{C}_{\mathbb{S_{Q}}}(\mathbb{C})$ is the centralizer of $\mathbb{C}$ [46].
-Note that $\left\langle\sigma|\tau\right\rangle=d^{|\sigma\tau^{-1}|}$ where $d$ is the local Hilbert space dimension and $|\sigma\tau^{-1}|$ denotes the number of cycles in permutation $\sigma\tau^{-1}$.
-In the $d\to\infty$ limit, $\left\langle\sigma|\tau\right\rangle$ becomes $d^{Q}\delta_{\sigma,\tau}$.
-This corresponds to a random field $Q!$-state Potts model, where the first term can be interpreted as a ferromagnetic coupling, the second term as two competing random fields, and the final term as the effect of temperature.
-The strength of the random fields is controlled by $q$ while temperature is controlled by $p$.
-Given prior work demonstrating phase transitions in random field Ising models and Potts models [50, 51, 52, 53, 54, 55, 56, 57, 58, 59], we anticipate that a phase transition would occur at finite $p$ and $q$ in quantum circuit which is directly characterized by $\overline{S(M|A)}$.
-For depolarizing and dephasing noises, the corresponding statistical-mechanics models are analogous as provided in Supplemental Material [46].
+In the absence of random fields, the model exhibits spin permutation symmetry [47].
+However, any non-zero net external field explicitly breaks this symmetry.
+As a result, the bulk spins become aligned with either $\mathbb{I}$ or $\mathbb{C}$, leading to the formation of domain walls due to the imposed boundary conditions.
+The free energy cost associated with these domain walls scales proportionally with the length of the subsystem, resulting in a volume-law scaling of the CEE.
+A transition from volume-law to area-law scaling becomes possible only when the net external field vanishes, restoring the spin permutation symmetry between $\mathbb{I}$ and $\mathbb{C}$.
+This requires that the QE operation is symmetric to the noise, along with equal rates for them.
+Specifically, unless otherwise indicated, we assume the symmetric scenario where $U_{SE}=U_{SA}$, $\rho_{E}=\rho_{A}$ and $q_{n}=q_{e}=q/2$ in the following.
+Given prior work demonstrating phase transitions in random field Ising models and Potts models with zero-field conditon [48, 49, 50, 51, 52, 53, 54, 55, 56, 57], we anticipate that a phase transition would occur at finite $p$ and $q$ in quantum circuit.
+At high temperatures (large $p$), the system enters a paramagnetic phase where domain walls are absent, resulting in area-law scaling.
+Conversely, at low temperatures (small $p$), the symmetry is spontaneously broken and domain walls form along the boundary, yielding volume-law scaling.
 Figure 3: Relationship between symmetries.
@@
 Since $\overline{S(M|A)}$, as opposed to $\overline{S(M)}$ in traditional setting, undergoes a phase transition, it is pertinent to determine under what conditions $\overline{S(M|A)}$ serves as a valid probe of entanglement or quantum information [58].
+This ensures that the observed phase transition constitutes a genuine entanglement phase transition with meaningful operational significance.
 It has been proven that if the system has information exchange (IE) symmetry in every realization, such that $S(M|A)=S(M|E)$, then $S(M|A)$ ensures entanglement between $M$ and $M^{c}\cup A$ [44].
@@
 The aAEE symmetry requires that the probabilities of these two realizations occurring are identical.
-Concretely, one can represent the circuit’s outcome as $\rho=\sum_{m}\left|\Psi_{m}\right\rangle\left\langle\Psi_{m}\right|$, and the aAEE symmetry necessitates that
+Concretely, one can represent the circuit’s outcome as $\rho=\sum_{m}\left|\Psi_{m}\right\rangle\left\langle\Psi_{m}\right|$ [59], and the aAEE symmetry necessitates that
 $\text{SWAP}^{\prime}\rho\text{SWAP}^{\prime\dagger}=\rho.$ (3)
-This is reminiscent of the recently proposed average (weak) symmetry in the study of mixed state order [61, 62, 63, 64, 65, 66].
-A subtlety in defining $\rho$ in this manner is that the number of qubits in $\left|\Psi_{m}\right\rangle$ may vary, but this can be simply addressed by adding unentangled qubits to states with fewer qubits.
-With aAEE symmetry and combining with Eq.
-(2), one would have
-$\sum_{m}\Lambda_{A}^{Q}$ $=\text{Tr}_{A\cup E}(\rho^{\otimes Q}\mathbb{C}^{A})=\text{Tr}_{A\cup E}\left(\text{SWAP}^{\prime}\rho^{\otimes Q}\right.$ (5)
-$\left.\text{SWAP}^{\prime\dagger}\mathbb{C}^{A}\right)=\text{Tr}_{A\cup E}(\rho^{\otimes Q}\mathbb{C}^{E})\equiv\sum_{m}\Lambda^{Q}_{E}.$
-This result implies that $\overline{S(M|A)}=\overline{S(M|E)}$, which we call the aIE symmetry. It is then straightforward to show that aIE symmetry results in,
-$\overline{S(M|A)}\geq 0,\quad\overline{S(M|A)}=\overline{S(M^{c}|A)},$
-$\overline{S(M|A)}=-\overline{S(M|A\cup M^{c})},$ (6)
-which also ensure that $\overline{S(M|A)}$ represents the entanglement between $M$ and $M^{c}\cup A$.
-Translating the aAEE symmetry into the concrete circuit model reveals that it requires $\rho_{E}=\rho_{A}$, $U_{SE}=U_{SA}$ and $q_{n}=q_{e}=q/2$.
-Thus, QE operation should be symmetric with the corresponding noise, and the operation should be applied at the same rate as the noise.
-It is noteworthy that aAEE symmetry condition $q_{n}=q_{e}$ aligns with the condition of zero net field in the statistical-mechanics model.
-Since every noise can be represented in the Stinespring form [45], QE operations can protect MIPT from various types of noise.
+This is reminiscent of the recently proposed average (weak) symmetry in the study of mixed state order [60, 61, 62, 63, 64, 65, 66]. With aAEE symmetry and combining with Eq. (2), one would have
+$\sum_{m}\Lambda_{A}^{Q}$ $=$ $\text{Tr}_{A\cup E}(\rho^{\otimes Q}\mathbb{C}^{A})=\text{Tr}_{A\cup E}\left(\text{SWAP}^{\prime}\rho^{\otimes Q}\text{SWAP}^{\prime\dagger}\right.$ (4)
+$\left.\mathbb{C}^{A}\right)=\text{Tr}_{A\cup E}(\rho^{\otimes Q}\mathbb{C}^{E})\equiv\sum_{m}\Lambda^{Q}_{E}.$
+This result implies that $\overline{S(M|A)}=\overline{S(M|E)}$, which we call the aIE symmetry. It also ensures that $\overline{S(M|A)}$ represents the entanglement between $M$ and $M^{c}\cup A$.
+Translating the aAEE symmetry into the concrete circuit model reveals that it actually coincides with the zero-field condition, which is necessary for satisfying Eq.
+(3).
+Therefore, the MIPT captured by CEE indeed represents the transition in the entanglement scaling law of the underlying quantum state.
+Moreover, since any noise can be represented in the Stinespring form [45], QE operations are capable of protecting MIPT from various types of noise.
 Figure 4: Numerical results for dephasing noise.
@@
 An immediate application of our scheme requires prior knowledge of the probability that a specific type of decoherence noise will occur in the experimental setting.
-Our scheme can be directly adapted to determine the noise rate as follows: apply a finite amount of measurement such that the state enters the area-law phase, then sweep through different QE operation rates $q_{e}$.
-In the thermodynamic limit, curves corresponding to different $p$ would intersect at the point $q_{n}=q_{e}$.
+Our scheme can be directly adapted to determine the noise rate as follows: apply a finite amount of measurement such that the state enters the area law phase, then sweep through different QE operation rates $q_{e}$.
+As long as the circuit is deep in the area law phase, the bipartite CEE should not change upon varying $p$, once the condition $q_{n}=q_{e}$ is met.
+Therefore, curves of CEE corresponding to different $p$ would intersect at the point $q_{n}=q_{e}$ in the thermodynamic limit.
+By performing finite-size scaling on the intersection point, we obtain an estimate of the noise rate $q_{n}$ in the system.
 A more detailed explanation and finite size numerical results are provided in [46].
@@
 Here the post-selection problem persists when the measurement rate is non-zero because determining the CEE of a single trajectory requires post-selecting on the measurement outcomes, a difficulty rooted in Born’s rule.
-Recent work has utilized various post-selection-free methods to probe the phase transition, such as using cross-entropy benchmarking [72] or providing bounds on the entanglement entropy [73, 74].
+Recent work has utilized various post-selection-free methods to probe the phase transition, such as using cross-entropy benchmarking [72, 73] or providing bounds on the entanglement entropy [74, 75].
 We believe these methods can also be extended to the conditional entropy in this work, and we leave this exploration for future work.
@@
 Kelly for valuable discussions.
-This work is supported by the National Key Research Program of China under Grant No.
-2019YFA0308404, the Natural Science Foundation of China through Grants No.
+This work is supported by the Natural Science Foundation of China through Grants No.
 12350404 and No.
 12174066, the Innovation Program for Quantum Science and Technology through Grant No.
-2021ZD0302600, the Science and Technology Commission of Shanghai Municipality under Grants No.
-23JC1400600 and No.
+2021ZD0302600, the National Key Research Program of China under Grant No.
+2019YFA0308404, the Science and Technology Commission of Shanghai Municipality under Grants No.
+23JC1400600, No.
+24LZ1400100 and No.
 2019SHZDZX01.